OpenStudy (jpsmarinho):

How to integrate this: 1/[(x+1)(x+2)] dx I tried integrate but With unsuccessful

OpenStudy (anonymous):

partial fractions will do it

OpenStudy (anonymous):

use Partial fractions.

OpenStudy (jpsmarinho):

Aah thanks! I don't know this in calculus

OpenStudy (anonymous):

\[\frac{1}{(x+1)(x+2)}=\frac{a}{x+1}+\frac{b}{x+2}\]

OpenStudy (jpsmarinho):

What is a and b ?constants?

OpenStudy (anonymous):

yes they are to be determined. are you familiar with Partial fractions ?

OpenStudy (jpsmarinho):

No =P. . And I need to learn this very quickly to a exam

OpenStudy (jpsmarinho):

Please, someone could solve for me the integral? It would a great help!

OpenStudy (anonymous):

continuing from where satellite left Multiply both sides with (x+1)(x+2) you will get 1=a(x+2)+b(x+1) put x=-1 in this you will get 1=a(-1+2)+b(0) OR a=1 now put x=-2 in 1=a(x+2)+b(x+1), you will get 1=a(0)+b(-2+1) b=-1 so partial fractions are \[\Large \frac{1}{(x+1)(x+2)}=\frac{1}{x+1}-\frac{1}{x+2}\] integrate right hand side \[\Large \int\limits_{}^{}\frac{1}{x+1}dx-\int\limits_{}^{}\frac{1}{x+2}dx\]

OpenStudy (anonymous):

@Fellowroot Plz do not provide whole solution.

OpenStudy (fellowroot):